ok the question is Refere to the right rectangular prism with the deminsions of L,W, and H. ON the prism is has the L=26 W=5 and the H=20. how do you find the Lateral Area and the Surface Area?
Answer by stanbon(75887)

Lateral Area does not include top and bottom
Front + Back = 2[5(20)]= 200 sq units
Left + Right = 2[26(20)]=1040 sq units
Lateral Area = 200 + 1040= 1240 sq units
Surface Area includes the top and bottom:
Top + Bottom = 2[5(26)]=260 sq units
Surface Area = 1240+260 = 1500 sq units
Cheers,
Stan H.